the sum of two numbers is one-hundred seventy-eight. The difference between the number is six. what is the product of the two numbers.
Answer by ankor@dixie-net.com(22740)

the sum of two numbers is one-hundred seventy-eight.
a + b = 178
:
The difference between the number is six.
a - b = 6
:
Use elimination here,
a + b = 178
a - b = 6
-----------adding eliminates b, find a
2a = 184
a = 184/2
a = 92
then
92 + b = 178
b = 178 - 92
b = 86
:
what is the product of the two number, I'll let you find the product